A supply officer distributes 80 business envelops among 13 employees, giving half the number to the female employees and the other half to the male employees. Each female employee received 3 more than each male. How many male employees are there? ( answer is 8 - show your solution )
Answer by stanbon(75887)

A supply officer distributes 80 business envelops among 13 employees, giving half the number to the female employees and the other half to the male employees. Each female employee received 3 more than each male. How many male employees are there? ( answer is 8 - show your solution )
------------------------
Let the number of males be "x"; the number of females be "13-x"::
------
# each male receives = 40/x
# each female receives = 40/(13-x)
------
Equation:
40/(13-x) - 40/x = 3
------------------------
40x - 40(13-x) = 3x(13-x)
------
80x - 40*13 = 39x - 3x^2
-------
3x^2 + 41x - 40*13 = 0
----
(x-8)(3x+65) = 0
Positive solution::
# of males = x = 8
--------------------------
Cheers,
Stan H.
------------------